# Towards Immersive Mixed Reality Street Play

### v1

@@ 5. Findings
[…]
@@ 5.1. Social Interactions
[…]
We observed that players quickly formed connections with each other during play, and the presence of an audience introduced performative elements. The game blurred the line between players and spectators, creating new opportunities for engagement in the public setting.
5.1.1. Collaboration and competition (31/65, behavior)
@@
In interviews, participants consistently favored the multiplayer aspect of the game, noting that competition and collaboration both strengthened social ties. They emphasized the unique excitement of playing together in person. One player explained, “If there are many people, and there is a certain sense of excitement and competition” (P4), suggesting that having more people involved made the experience more thrilling. Others simply enjoyed this way to spent time with friends and family, remarking that “Anyway, it’s interesting to have a new way to play with people” (P27). As another participant put it, “Probably the interesting part is not the magic, but the person on the opposite side” (P10), highlighting that the real fun came from interacting with other people rather than the technology itself.
5.1.2. Showmanship and performance (25/65, behavior)
Because IMRSP unfolds in public and is highly visible, it often encourages participants to show off and perform. The game provides a unique opportunity for adults to engage in playful, out-of-the-ordinary behaviors in a public setting, which draws attention from bystanders and creates a shared sense of excitement. In the presence of an audience, players sometimes embellished their actions with dramatic flourishes purely for entertainment, even if these gestures did not improve their chances of winning. We saw participants adding theatrical body movements to their play, contributing to the overall spectacle for themselves and onlookers. Some examples of this playful showmanship included:
@@ 5.1.2. Showmanship and performance (25/65, behavior)
One particularly popular bit of showmanship was the ritual of clicking wands together before a match. This gesture—featured in the game’s tutorial video—quickly became a customary way for players to bond; many would click their wands together both before and after the game for photos or simply as a friendly greeting (P1–2, P11–12, P15–16, P30–32, P50–53). (see Figure 7).
@@
Figure 7. Clicking the wands before or after the play becomes popular among participants.
5.1.3. Spectating and honey pot effect (41/65, behavior)
Figure 8. (a) Passersby being attracted by players’ body movements, and then (b) approached the spectator view and took pictures.
When played in public spaces, MR street play invokes the interest of passersby. MR street play is fully observable to the public, and passersby can show strong interest in the game.
People may stop to spectate for various reasons, including being attracted by the player’s body movements, the spectators’ view, or being attracted by other observing people (see Figure 8).
Observed behaviors include
  • Took photos or videos of the players during the game (12/65)
  • Recorded the spectator view screen to capture the mixed-reality action (7/65)
  • Cheered for the players (25/65)
  • Shared photos or videos of the scene with friends or on social media (7/65)
  • Chatted with other bystanders about the game (34/65)
  • Volunteered to join the next game session themselves (41/65).
Reviews highlight the importance of the spectator mode, with one reviewer noting, ”Spectator mode was also an incredible idea” (L14). Participants also expressed increased curiosity and desire to join the game after seeing the spectator view, with comments like ”After I see the spectator view…I really want to join” (P10) and ”This is sick I want to play too” (P31).
The spectator view also adds a social and fun element to the game, as one participant mentioned, ”If they have the spectator view, the bystanders can also comment on the game when the two of us are dueling, this is more fun” (P3). However, not having access to the MR view can be detrimental to the experience, ”If he can’t see MR at all, he’s actually an obstacle in my game for me” (P9).
5.1.4. Inclusive Multiplayer with Hand-Held Mobile Device (8/65, behavior)
The controller role on hand-held devices turns out to be more popular than the HMD player roles, despite many participants not having tried the game in HMD before. When the researchers showcased the dragon hunting game, all players and spectators eagerly volunteered to be the dragon controller, even if they had never used the HMD or played the wizard duel game yet. One participant stated ”Being the boss behind the scenes sounds totally sick” and ”I wanna blast my friends with some fire!” (P32). Another shared their surprise at the addition of the controller role, stating, ”Fight the Dragon gave me a new AR experience, because I used to think only of two-player battles, and I didn’t expect that there would be a third machine to control the character” (P18).
@@ 5.1.5. Appreciation of AR over VR (8/63)
People expressed a strong preference for augmented reality (AR) experiences over virtual reality (VR) when engaging with the MR street play MOFA, both on-site and online. The ability to interact with friends and other facilities in the real world, rather than fully immersing in a virtual environment, was a key factor in their enjoyment and interest in the game. As one participant stated, ”I am way more interested in a digital world being overlaid on the real world - augmenting rather than escaping” (L22). Other participants highlighted the limitations of VR, mentioning, ”You can’t even pick up a beer while playing VR”; ”You can’t see around you in Quest unless you tap on the side to see it to the past as through me personally, I like to be immersed in the game while also being able to see what’s around” (P19).
@@ 5.2. Social Concerns
[…]
While the social benefits of MR street play were evident, our study also uncovered significant social concerns and friction points, particularly regarding how this novel activity is perceived within broader society. These concerns were more frequently and vividly expressed in the online reviews, though some on-site participants acknowledged them when prompted. We detail these issues below.
@@ 5.2.1. Social awkwardness and perceived impropriety (13/63)
13 reviewers, all from the online forum data collection, expressed feelings of awkwardness as their main impression of the game, with some feeling offended by the prospect of engaging in immersive MR street play. The reviewers questioned the social acceptance level of wearing the device in public, particularly regarding the potential impact on non-participants:
”I’m not sure if I’d like to walk around the streets wearing them though - would be incredibly annoying to other people.” (L1)
”Disregard for culture and obvious social media reactions.” (L19)
One reviewer expressed empathy for passersby who might be unintentionally captured in the middle of others’ game space, noting that they ”don’t look too pleased” (L26). The impropriety of engaging the game in certain contexts was also highlighted, with one reviewer commenting, ”It seems pretty impolite to be LARPing in a busy train station” (L10).
The rest comments expressed more straight-forward emotions towards this phenomenon as ”playing complete idiots in public places” (L2), or ”dork” or “dorky” to using HMD on streets (4/63), and refer the MR street players as ”being a nuisance to everyone” (L27).
The aesthetics and form factor of the MR devices were also a point of concern, with one reviewer comparing the experience to strapping ”a heavy awkward thing to the front of my face” (L5).
These responses reflect explicit social disapproval of public immersive MR games. It underscores the challenges of overcoming social awkwardness and perceived impropriety. Addressing these concerns will be crucial in fostering widespread adoption and acceptance of this emerging form of entertainment in public spaces.
@@
Regarding the on-site participants, only 1 participant showed slight hesitation and discomfort when first introduced to the headset in public, who later accepted it and tried the game (P20). Other on-site participants didn’t express feelings of awkwardness or initiate this topic in interviews until the researchers asked about their feelings of social awkwardness in the interview.
5.2.2. Establishing social norms
@@
While some users expressed concerns about the social awkwardness and perceived impropriety of using MR technology in public spaces, others held a more optimistic view. They believed that the adoption of new technologies has historically led to changes in social norms, and MR street play is likely to follow a similar path (7/63).
@@
Several users drew comparisons to the normalization of other technologies and behaviors that were once considered socially unacceptable or embarrassing. One participant noted that talking with headphones is now socially acceptable, despite the initial awkwardness of ”talking to the air” (P20). Another user pointed out that glasses, umbrellas, and bikes were all ”embarrassing at first until it becomes normalized” (L23). Some participants expressed confidence that technological evolution would lead to reduced headset sizes, drawing parallels to the historical miniaturization of mobile phones from their initial ”brick-like” dimensions to current compact forms that have gained widespread social acceptance (L22).
@@
The widespread adoption of Pokémon Go was also cited as an example of how a seemingly ”dorky” behavior can become socially acceptable when enough people engage in it. Users agreed that people would wear headsets outside if the experience is compelling enough, just as there was ”a plague of arms-length phone-holders in every park for a year” during the height of Pokémon Go’s popularity (L25). The social acceptability of the behavior was further reinforced when enjoyed in the company of others, as one user noted, ”If you and a friend are doing this together, you’re no longer a dork” (L31). These comments suggest that there are people believing that as MR technology becomes more compelling and widely adopted, the social norms surrounding its use in public spaces are likely to evolve.
@@
However, some others pointed out that the social norm surrounding MR technology in public settings has yet to be established (4/63). One user highlighted the difference between the social acceptability of looking at a phone and engaging in MR street play, stating, ”Waving a stick around and shouting fireball while wearing a goofy headset is, at present, still unacceptable” (L8). Others suggested that the widespread adoption would require the integration of a popular AR game like Pokémon GO (L11).
@@
The form factor of the device was identified as a crucial factor in determining the social acceptability of MR street play. As one user pointed out, ”This is not the form factor that will change the paradigm. The smartphone was…Whoever gets the AR glasses that look like glasses out first wins the game” (L14). Another user doubted the establishment of a new social norm surrounding MR in the reply as “This will only happen if the devices start to become indistinguishable from a normal pair of glasses” (L12).
5.2.3. Gradients of Social Awkwardness
Our findings revealed a clear hierarchy in social acceptance levels, where multiplayer scenarios demonstrated lower levels of social awkwardness compared to single-player experiences. Furthermore, active players consistently reported less social discomfort than bystanders.
@@ 5.2.3. Gradients of Social Awkwardness
First, players in multiplayer games are less likely to feel awkward compared to single-player experiences. They feel more comfortable when the MR world is shared with another person, knowing they are not alone in MR while being observable to passersby. From the passersby’s perspective, watching the players’ interactive body movements gives them hints about the content of the MR game. An online reviewer commented, ”If you and a friend are doing this together, you’re no longer a dork. You’re going your own way” (L31). When questioned about awkward levels in public settings, all on-site participants reported lower awkward feelings during paired games than single-player games. ”2 people playing the game out there is fully acceptable” (P1-12). One of them stated,
”It’s no longer embarrassing if it takes two doing embarrassing things together.” (P12)
@@
Secondly, participants are also less likely to feel embarrassed than spectators or passersby. When participants recalled their feelings before they tried the game, they described as bystanders they would feel confused and anxious as not knowing what was actually happening in the MR experience. In contrast, after the game, even if playing in single-player mode and behaving abnormally, many claimed they did not feel embarrassed but were eager to explain the content and gameplay to others. As one participant put it, ”It’s not embarrassing to play, once you start it’s someone else’s concern to be embarrassed” (P6).
However, opinions vary on whether playing with strangers is more awkward than playing alone in public spaces. The spectator view can help alleviate awkwardness but does not fully change people’s willingness. More introverted individuals tend to stick to close friends, while some extroverted people prefer playing with others over playing alone in public.
5.2.4. Safety and security concern (7/63)
@@
Many people also expressed practical concerns about the potential safety issues that immersive MR street play can result in. There is a common concern about tread accidents due to the players being too immersed in the MR game and ignoring the context in reality (5/63). One person commented online about the MOFA videos: ”I do worry that one of my kids is going to get trampled by an adult who can’t see because he’s too busy slaying a dragon.” (15)
The other major concern is about the HMD device being stolen (2/63). People worried that putting an iPhone in the headset would bring a ”significant risk of phone theft in major cities” (L25).
5.2.5. Divergence of Online and Onsite Review
The rate of negative reviews online is considerably greater than that of onsite reviews. Online reviews demonstrate a notable imbalance, with a positive-to-negative feedback ratio of 19:20, with some posting both interests and concerns. In contrast, on-site feedbacks reflect a much higher ratio of 70:12 in favor of positive comments.
@@ 5.2.5. Divergence of Online and Onsite Review
Moreover, the content of negative feedback also differs in the 2 data collections. On-site negative comments are often about specific technical limitations, including the field of view limitations and anchoring issues (10/32), as described in the Gameplay Experience section. These critiques are typically framed in a manner that suggests a desire for improvement, with a more subdued and neutral tone. Most on-site participants (31/32) didn’t express concerns about social acceptance until the researchers prompted them on this topic.
@@
In contrast, online feedback is significantly more critical and confrontational. Online reviewers frequently raise pointed questions challenging the social acceptance of the game (13/31), potential safety risks associated with gameplay (3/31), MR headset form factors (3/31), and even the overall entertainment value of the content (2/31).
@@ 5.3. Gameplay Experience
Onsite participants unanimously liked the game experience, which was described as ”fun”, ”amazing”, ”incredible”, and “magical” (31/63). They particularly liked that the game was intuitive to understand and operate and that they could play and communicate with other players. However, some players also expressed frustration due to technological limitations.

### v2

@@ 5. Findings
In this section, we present our findings, which are organized into four key themes that emerged from the data analysis. These themes capture different facets of the IMRSP experience and provide a framework for understanding our findings. We complement our qualitative insights with behavioral coding, using the notation $x/y$ to show that $x$ out of $y$ participants demonstrated a specific behavior or attitude. Specifically, $y=65$ for on-site participants’ behaviors (P1–P65); $y=63$ for participant and respondent attitudes (P1–P32, L1–L31). We link quotes with personal IDs to indicate whether the insights come from on-site participants (e.g., P1) or online respondents (e.g., L1).
@@ 5.1. Social Interactions
We observed that players quickly formed connections with each other during play, and the presence of an audience introduced performative elements. The game blurred the line between players and spectators, creating new opportunities for engagement in a public setting.
5.1.1. Collaboration and competition (31/65)
@@
In interviews, participants consistently favored the multiplayer aspect of the game, noting that competition and collaboration both strengthened social ties. They emphasized the unique excitement of playing together in person. One player explained, “If there are many people, and there is a certain sense of excitement and competition” (P4), suggesting that having more people involved made the experience more thrilling. Others simply enjoyed this way to spend time with friends and family, remarking that “Anyway, it’s interesting to have a new way to play with people” (P27). As another participant put it, “Probably the interesting part is not the magic, but the person on the opposite side” (P10), highlighting that the real fun came from interacting with other people rather than the technology itself.
5.1.2. Showmanship and performance (25/65)
Because IMRSP unfolds in public and is highly visible, it often encourages participants to show off and perform. The game provides a unique opportunity for adults to engage in playful, out-of-the-ordinary behaviors in a public setting, which draws attention from bystanders and creates a shared sense of excitement. In the presence of an audience, players sometimes embellished their actions with dramatic flourishes purely for entertainment, even if these gestures did not improve their chances of winning. We observed participants incorporating theatrical body movements into their play, thereby contributing to the overall spectacle for themselves and onlookers. Some examples of this playful showmanship include:
@@ 5.1.2. Showmanship and performance (25/65)
One particularly popular instance of showmanship was the ritual of clicking wands together before a match (see Fig. 8). This gesture—featured in the game’s tutorial video—quickly became a customary way for players to bond; many would click their wands together both before and after the game for photos or simply as a friendly greeting (P1, P2, P11, P12, P15, P16, P30–P32, P50–P53).
@@
Figure 8. Players develop a ritual of playful showmanship: they click wands together with teammates or opponents before or after play, establishing a popular game tradition that resembles the wand interactions of wizards in Harry Potter movies.
5.1.3. Spectating and honeypot effect (41/65)
Figure 9. Honeypot effect observed: (a) Passersby were attracted by players’ body movements, then (b) approached the researcher holding the spectator view of the game, watched the augmented game view, and took pictures of it.
When played in public spaces, MR street play invokes strong interest from passersby as it is fully observable, and arguably hard not to notice. People stop to spectate for various reasons, including being attracted by the player’s body movements, the spectators’ view, or being attracted by other observing people (see Fig. 9). Typical observed behaviors of the spectators include:
  • Taking photos or videos of the players during the game (12/65)
  • Recording the spectator view screen to capture the mixed-reality action (7/65)
  • Cheering for the players (25/65)
  • Sharing photos or videos of the scene with friends or on social media (7/65)
  • Chatting with other bystanders about the game (34/65)
  • Volunteering to join the next game session themselves (41/65)
Participants expressed increased curiosity and desire to join the game after seeing the spectator view, with comments like “After I see the spectator view…, I really want to join” (P10) and “This is sick! I want to play too” (P31). In addition, the online reviews highlight the importance of the spectator mode, with one comment noting, “spectator mode was also an incredible idea” (L14).
The spectator view was also found to add a social and enjoyable element to the game, as one participant (P3) mentioned, “If they have the spectator view, the bystanders can also comment on the game when the two of us are dueling, this is more fun.” However, not having access to the MR view can be detrimental to the experience, as P9 noted, “If he can’t see MR at all, he’s actually an obstacle in my game for me.
5.1.4. Inclusive multiplayer with handheld mobile device (8/65)
The controller role on handheld devices turns out to be more popular than the MRHMD player roles, despite many participants not having tried the game in HMD before. When the researchers showcased the dragon hunting game, all players and spectators eagerly volunteered to be the dragon controller, even if they had never used the HMD or played the wizard duel game yet. One participant stated “Being the boss behind the scenes sounds totally sick” and “I wanna blast my friends with some fire!” (P32). Another shared their surprise at the addition of the controller role, stating, “Fight the Dragon gave me a new AR experience, because I used to think only of two-player battles, and I didn’t expect that there would be a third machine to control the character” (P18).
@@ 5.1.5. Appreciation of AR over VR (8/63)
People expressed a strong preference for augmented reality experiences over virtual reality, both on-site and online. The ability to interact with friends and other facilities in the real world, rather than fully immersing in a virtual environment, was a key factor in their enjoyment and interest in the game. As one online comment stated, “I am way more interested in a digital world being overlaid on the real world—augmenting rather than escaping” (L22). Other participants highlighted the limitations of VR, mentioning, “You can’t even pick up a beer while playing VR” (L18); “You can’t see around you in Quest unless you tap on the side… personally, I like to be immersed in the game while also being able to see what’s around” (P19).
@@ 5.2. Social Concerns
While the social benefits of MR street play were evident, our study also uncovered significant social concerns and friction points, particularly regarding how this novel activity is perceived within the broader society. These concerns were more frequently and vividly expressed in the online reviews, though some on-site participants acknowledged them when prompted. We detail these issues below.
@@ 5.2.1. Social awkwardness and perceived impropriety (13/63)
Thirteen respondents (13/63), all from the online forum data collection, expressed feelings of awkwardness as their main impression of the game, with some feeling offended by the prospect of engaging in immersive MR street play. These comments questioned the social acceptance level of wearing the device in public, particularly regarding the potential impact on non-participants. For example, “I’m not sure if I’d like to walk around the streets wearing them though—it would likely be incredibly annoying to other people” (L1). L19 criticized the play as “disregard for culture and obvious social media reactions.” One comment expressed empathy for passersby who might be unintentionally captured in the middle of others’ game space, noting that they “don’t look too pleased” (L26). The impropriety of engaging in the game in certain contexts was also highlighted, with L10 commenting, “It seems pretty impolite to be LARPing in a busy train station”. The remaining comments expressed more straightforward emotions towards this phenomenon, accusing the MR street players of being “a nuisance to everyone” (L27), and finding it to be “complete idiots to play in public places” (L2), and “dorky” to use HMD on streets. In addition, the aesthetics and form factor of the MR devices were also a point of concern, with one comment comparing the experience to strapping “a heavy, awkward thing to the front of my face” (L5). These responses reflect explicit social disapproval of public immersive MR games. They underscore the challenges of overcoming social awkwardness and perceived impropriety. It is crucial to address these concerns to foster widespread adoption and acceptance of this emerging form of entertainment in public spaces.
@@
Regarding the on-site participants, only one participant showed slight hesitation and discomfort when first introduced to the headset in public, who later accepted it and tried the game (P20). Other on-site participants did not express feelings of awkwardness or initiate this topic in interviews until the researchers asked about their feelings of social awkwardness in the interview.
5.2.2. Establishing social norms (7/63)
@@
While some users expressed concerns about the social awkwardness and perceived impropriety of using MR technology in public spaces, others held a more optimistic view. They believed that the adoption of new technologies has historically led to changes in social norms, and MR street play is likely to follow a similar path.
@@
Several users drew comparisons to the normalization of other technologies and behaviors that were once considered socially unacceptable or embarrassing. One participant noted that talking with headphones is now socially acceptable, despite the initial awkwardness of “talking to the air” (P20). Another user pointed out that glasses, umbrellas, and bikes were all “embarrassing at first until it becomes normalized” (L23). Some participants expressed confidence that technological evolution would lead to reduced headset sizes, drawing parallels to the historical miniaturization of mobile phones from their initial “brick-like” dimensions to current compact forms that have gained widespread social acceptance (L12).
@@
The widespread adoption of Pokémon GO was also cited as an example of how a seemingly “dorky” behavior can become socially acceptable when enough people engage in it. Users agreed that people would wear headsets outside if the experience is compelling enough, just as there was “a plague of arms-length phone-holders in every park for a year” during the height of Pokémon GO’s popularity (L25). The social acceptability of the behavior was further reinforced when enjoyed in the company of others, as one user noted, “If you and a friend are doing this together, you’re no longer a dork” (L31). These comments suggest that there are people who believe that as MR technology becomes more compelling and widely adopted, the social norms surrounding its use in public spaces are likely to evolve.
@@
However, some others pointed out that the social norm surrounding MR technology in public settings has yet to be established (4/63). One user highlighted the difference between the social acceptability of looking at a phone and engaging in MR street play, stating, “Waving a stick around and shouting fireball while wearing a goofy headset is, at present, still unacceptable” (L8). Others suggested that the widespread adoption would require the integration of a popular AR game like Pokémon GO (L11).
@@
The form factor of the headset device was identified as a crucial factor in determining the social acceptability of MR street play. As one user pointed out, “This (HoloKit with iPhone) is (yet) not the form factor that will change the paradigm, the smartphone is. Whoever gets the AR glasses that look like glasses out first wins the game” (L14). Another user doubted the establishment of a new social norm surrounding MR in the reply, stating that “This will only happen if the devices start to become indistinguishable from a normal pair of glasses” (L12).
5.2.3. Gradient of Social Awkwardness
Our findings revealed that a gradient of social comfort levels emerged across different modes of engagement: multiplayer play felt the most natural and least awkward, while solo play introduced some discomfort. Watching others play without access to the spectator view was described as the most awkward experience.
@@ 5.2.3. Gradient of Social Awkwardness
First, players in multiplayer games are less likely to feel awkward compared to single-player experiences. They feel more comfortable when the MR world is shared with another person, knowing they are not alone in MR while being observable to passersby. From the passersby’s perspective, watching the players’ interactive body movements gives them hints about the content of the MR game. When questioned about awkward levels in public settings, all on-site participants reported a lower sense of awkward feelings during paired games than single-player games. P1-P12 agreed that two people playing the game out there is fully acceptable. One of them stated, “It’s no longer embarrassing if it takes two doing embarrassing things together” (P12).
@@
Secondly, participants are also less likely to feel embarrassed than spectators or passersby. When participants recalled their feelings before they tried the game, they described feeling confused and anxious, as they did not know what was happening in the MR experience. In contrast, after the game, even if playing in single-player mode and behaving abnormally, many claimed they did not feel embarrassed but were eager to explain the content and gameplay to others. As one participant put it, “It’s not embarrassing to play, once you start it’s someone else’s concern to be embarrassed” (P6).
5.2.4. Safety and security concerns (7/63)
@@
Many people expressed practical concerns about the potential safety issues that immersive MR street play can result in (5/63). There is a common concern about collision accidents due to the players being too immersed in the MR game and ignoring the real-world context among online respondents. One person commented about the MOFA videos, “I do worry that one of my kids is going to get trampled by an adult who can’t see because he’s too busy slaying a dragon” (L17). Another major concern is about the HMD device being stolen (2/63). People worried that putting an iPhone in the headset would bring a “significant risk of phone theft in major cities” (L24).
5.2.5. Divergence of online and on-site attitudes
Our analysis showed that the sentiment of online comments was more negative than that of on-site participants.
@@ 5.2.5. Divergence of online and on-site attitudes
Moreover, the content of negative feedback also differs. On-site negative comments are often about specific technical limitations, including anchoring issues, the field-of-view limitations, and tracking latency (11/32, see 5.3.4). These critiques are typically framed in a manner that suggests a desire for improvement, with a more subdued and neutral tone. Most on-site players (31/32) did not express concerns about social acceptance until the researchers prompted them on this topic.
@@
In contrast, online feedback is significantly more critical and confrontational. Online comments frequently raise pointed questions challenging the social acceptance of the game (13/31), potential safety risks associated with gameplay (3/31), MR headset form factors (3/31), and the overall entertainment value of the content (2/31).
@@ 5.3. Gameplay Experience
On-site participants generally described their game experiences as “fun”, “amazing”, “incredible”, and “magical” (31/63). They particularly appreciated that the game was intuitive to understand and operate, and that they could play and communicate with other players. However, some players also expressed frustration due to technological limitations, such as heavy headsets that caused discomfort on their noses (5/63).